A rectangular swimming pool is three times as long as it is wide and is surrounded by a deck 2.5m wide. Find the dimensions of the pool if the area of the deck is 265 meters squared.
Answer by ankor@dixie-net.com(22740)

A rectangular swimming pool is three times as long as it is wide and is surrounded by a deck 2.5m wide.
Find the dimensions of the pool if the area of the deck is 265 meters squared.
:
let x = the width of the pool
then
3x = the length of the pool
:
The overall length and with including the walkway a 2.5 wide deck will add 5' to the dimensions of the pool therefore
(x+5) by (3x+5) = the overall dimensions
:
Overall area - pool area = deck area
(x+5)(3x+5) - 3x*x = 265
3x^2 + 5x + 15x + 25 - 3x^2 = 265
20x = 264 - 25
20x = 240
x = 240/20
x = 12 meters is the width of the pool
and
3(12) = 36 meters is the length
z:
:
Check this find the area
overall area: 17 * 41 = 687
pool only area: 12*36 = 432
-----------------------------
deck area only: 265